OpenStudy (anonymous):

find the domain of the function f(x)= sqrtx/x-1 find the domain of the function f(x)= sqrtx/x-1 @Mathematics

OpenStudy (across):

x can't obviously be 1 and x can't obviously be less than zero.

OpenStudy (across):

\[\left \{ x\in\mathbb{R}:x\geq0 \wedge x \neq 1 \right \}\]

OpenStudy (anonymous):

\[\Large x \in [0,1) \cup (1,\infty )\]